OpenStudy (anonymous):

The floor of a rectangular room will be tiled with square tiles that measure 3 inches on each side. The room is 109 inches long and 103 inches wide. How many tiles, including parts of tiles, are needed to complete the job?

OpenStudy (anonymous):

A. 1,248 tiles B. 1,247 tiles C. 1,224 tiles D. 1,295 tiles

OpenStudy (anonymous):

you have to find the square feet first of the tile and then of the floor

OpenStudy (anonymous):

you could do it like this. 103 / 3 = 34.33 tiles for covering the wide side 109 / 3 = 36.33 tiles for covering the long side. now multiply 34.33 * 36.33 = 1247.44